A cannon sits atop a hill that is 10.5 m above the field below. The cannon shoots a shell with a velocity of 67 m/s at an angle of 54o. How far from the cannon will the shell hit the ground?
1
Expert's answer
2012-11-05T09:56:24-0500
Vertical velocity of a cannonis v*sin54o. Time, when it reaches the top of it's trajectory is
t=(v*sin54o)/g=5.53s. The height it's gonna reach is
H=H0+vtsin54o-gt^2/2=160.25m. Then it starting to fall freely with acceleration
of g. Time it's need to fall is t=sqrt(2H/g)=5.72s. So, the time it was in the
air is 5.53+5.72=11.25s. Horizontal velocity is v*cos54o. Distance it will cover
until it falls S=11.25*v*cos54o=260.41m.